A Chemist has m gm of salt water that is m% salty.How many grams of salt must be add to make a solution that is 2m % salty?
Answer by Alan3354(69443)

g = # of grams
m = concentration = %/100
---------
Amount of salt = g*m
w = amount of water
------------
g = s + w
m = s/(s+w) = s/g
Add x amount of salt
---------
2m = (s+x)/(g + x) = 2s/g
(s+x)/(g + x) = 2s/g
g(s+x) = 2s(g+x)
gs+gx = 2gs + 2sx
gx = gs + 2sx
gx - 2sx = gs
x(g-2s) = gs
x = gs/(g-2s)
s = mg
x = mg^2/(g - 2mg)
x = mg/(1 - 2m)